OpenStudy (wondermath):

If 4^x-4^x-1=24, what is the value of (2x)^x?

OpenStudy (wondermath):

let me rewrite it: 4^x-4^(x-1)=24

OpenStudy (wondermath):

can't get the answer D:

OpenStudy (ash2326):

4^x(1-4^-1)24 4^x(1-1/4)=24 4^x=24*4/3 4^x=32 taking log 2 on both sides log 4^x=log 32 . log is log to the base 2 xlog4=log32 x*2=5 x=2.5 or 5/2 we have to find (2x)^x (2*5./2)^(5/2) (5)^5/2 or root (5^5) or 25root(5)

OpenStudy (wondermath):

Thank you so much :D